OpenStudy (anonymous):

Electrical power (P) is measured in watts using the formula P = V · I where V is the voltage in volts and I is the current in amperes. If there are 60 volts running to a light bulb of 78 watts, how much current is flowing through the light bulb? A- 1.3 amperes B- 138 amperes C- 4,680 amperes D-18 amperes

OpenStudy (pottersheep):

so v = 60 p = 78 we have to find I

OpenStudy (pottersheep):

P = V*I so I = P/V

OpenStudy (pottersheep):

I = 78/60. I don't have a calculator with me but I guess that's 1.3 amperes, so A!

OpenStudy (anonymous):

plzhelp me